# Comparison of the Migdal transition probabilities in electron-atom inelastic cross sections

### v3

@@ 5 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The errors are $\sim 20\ \%$ for Ar at 10 keV and $\sim 60\ \%$ for Xe compared to the database value shown in Fig. 1. For Xe, the database value is limited and the error will be improved for higher energy because the plane wave and the Coulomb wave approximation used for $u$-channel processes becomes better for high energy. Also, the approximation is not good for excitation cross sections because the excited states are approximated to the plane wave states. Therefore, the expected errors for excitation are not reliable. The differences of threshold energy comes from the error of calculated bound energy. Actually, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ error with experimental data Ibe:2017yqa . Note that this result does not compare the transition probability from the inner shell electrons because the probability is dominated by the outermost electrons.

### v5

@@ 6 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The differences are $\sim 30\ \%$ for Ar at 10 keV and $\sim 80\ \%$ for Xe compared with the database value shown in Fig. 1. For Xe, the database value is limited and the differences may be improved for higher energy because the plane wave and the Coulomb wave approximation used for $u$-channel processes becomes better for high energy. Also, the approximation is not good for excitation cross sections because the excited states are approximated to the plane wave states. Therefore, the expected differences for excitation are not reliable. The differences of the low energy threshold come from the kinematical constraint, $E_{{\rm kin},I}>2\Delta E$. Also, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ differences with experimental data Ibe:2017yqa .